# Planet formation regulated by galactic-scale interstellar turbulence

### v1

@@ 0.20 Caveats for the ISM accretion rate:
In Extended Data Figure 8, we show the results of these calculations for a solar mass star in an ISM of density $\rho_{\mathrm{gas}}=100\,M_{\odot}$ pc-3. We find that capture rates are typically maximised for cloud sizes $R_{\mathrm{cloud}}\sim 0.1{-}1R_{\mathrm{BHL}}$. Assuming $R_{\mathrm{acc}}=R_{\mathrm{BHL}}$ and substituting in $\Delta v_{\mathrm{gas}}=1$ km s-1 for a solar mass star into equation 9 yields $\dot{M}_{\mathrm{BHL}}\approx 2.4\times 10^{-8}\,M_{\odot}$ yr-1. Inferring the comparable cloud capture accretion rate from Extended Data Figure 8 yields a very similar rate. In general, we expect that tidal cloud capture may play a significant role in disc assembly, alongside BHL accretion. This may increase the rate of interstellar gas capture by a factor of order unity, and may be the subject of future numerical experiment.
Face-on accretion: A third mechanism for ISM accretion is face-on disc accretion, where an existing disc produces a large physical cross-section.[88, 66] In this case, the ram pressure between a pre-existing disc and the ISM mediates the accretion process. Face-on accretion may be significant when a star moves through a relatively high velocity medium. In this case, the disc shrinks as a result of deposition of low angular momentum material onto the disc surface.[66] There is no steady state for this process unless material can be constantly replenished. This replenishment of outer disc material may be viscous, but then the quantity of material accreted is limited by the initial disc mass. Replenishment via BHL or cloud capture would keep the disc at a radius $R_{\mathrm{disc}}\sim R_{\mathrm{BHL}}$, and would therefore result in a similar accretion rate to BHL accretion.[88] However, this may significantly enhance accretion rates when the BHL radius decreases; then $R_{\mathrm{disc}}>R_{\mathrm{BHL}}$. Face-on accretion would then temporarily dominate the overall rate, shrinking the disc to maintain a balance between shrinkage and replenishment. Overall, we may expect some moderate enhancement to the rate of capture and stellar accretion during short phases, although this would not be expected to greatly change the overall disc evolution.
0.21 Consequences of ISM regulated discs:
@@
  • Disc sub-structures: A diverse range of substructures have been observed in protoplanetary discs.[48] We have demonstrated that the disc is undergoing constant perturbation by in-falling material. We would therefore expect some discs to exhibit structure due to this process.[17] For example, by accreting gas from the ISM with a misaligned angular momentum vector we expect disc warping,[135] which may also produce observable shadows.[136, 119] Both external structures and disc warping are evident in some discs,[9] and it seems probable that they originate from late stage in-fall at least in some instances.
@@
  • Accretion outbursts and short accretion times: FU Orionis events are short accretion bursts that last $\sim 100$ years,[137] named for the archetypal example.[138, 139] In-fall events are among the earliest explanations for these events. There remain several other possible mechanisms such as stellar encounters[140, 141] and thermal instability,[142, 143, 144] although extended nebulosity surrounding several stars undergoing outburst strongly supports the role of in-fall in at least some cases.[145, 146] Outbursting star-disc systems may be more broadly categorised as a subset of systems with short accretions time-scales.[42, 44] Among such discs, a lifetime problem appears particularly pronounced for discs around some Herbig stars.[147] The assumptions of our model mean that we cannot directly explain these short accretion time-scales; we impose this time-scale for our model. However, given a fixed time-scale $\tau_{\mathrm{acc}}$, discs with short $\tau_{\mathrm{acc}}$ may be episodically replenished in our model. Thus, to a degree the inferred accretion time-scale is decoupled from the expected frequency of observed discs, which no longer needs to be $\sim\tau_{\mathrm{acc}}/\tau_{\mathrm{disc}}$, for disc age $\tau_{\mathrm{disc}}$. Instead, the frequency of discs with a given $\tau_{\mathrm{acc}}$ would be dictated by the frequency of in-fall events that lead to rapid mass gain. Indeed, in our model a small number of discs with very short accretion time-scales ($\sim 0.01$ Myr) are retained for $\sim 3$ Myr (Extended Data Figure 4). Whether the in-fall itself can also produce a short-lived dip in the accretion time-scale (i.e. an outburst) depends on the turbulence imparted onto the disc, and may depend on the rate or physics of interstellar gas accretion (as discussed above).
@@
  • Long-lived discs: Stochastic replenishment of discs represents a simple explanation for so-called ‘Peter Pan’ discs, which are (low mass) stars that are tens of Myr old but still exhibit a gas-rich accretion disc.[148, 149, 108] It is not clear if Peter Pan discs are exclusively around low mass stars, or whether finding discs around M dwarf stars is just down to sampling of the mass function. Speculatively, this phenomenon may be related to the fact that individual massive discs appear across a number of regions of different ages, even where the majority of discs have been depleted.[150] Late stage in-fall seems a viable way to explain some of these long-lived discs. This further motivates deep observational programs to characterise the super-disc scale gas surrounding protoplanetary discs.
@@
  • Planet properties: Possibly the most exciting consequence of protoplanetary discs that are regulated by BHL accretion would be the potential consequences for planets. Multiple studies have shown that mature planet properties are dependent on stellar kinematics.[151, 152] Some of these differences could be related to covariance with stellar properties, such as age and metallicity.[153] However, it is not clear that all correlations can be explained by stellar properties alone.[154, 155, 156] If discs are constantly replenished by the ISM, stars that form at high velocity relative to the local rest frame should accrete less mass from the ISM after the dispersal of their formation cloud. This represents a direct theoretical link between planet formation and stellar kinematics.

### v2

@@ 4.1.2 Cloudlet capture
[…]
In Figure 13, we show the results of these calculations for a solar mass star in an ISM of density $\rho_{\mathrm{gas}}=100\,M_{\odot}$ pc-3. We find that capture rates are typically maximised for cloud sizes $R_{\mathrm{cloud}}\sim 0.1{-}1R_{\mathrm{BHL}}$. Assuming $R_{\mathrm{acc}}=R_{\mathrm{BHL}}$ and substituting in $\Delta v_{\mathrm{gas}}=1$ km s-1 for a solar mass star into equation 7 yields $\dot{M}_{\mathrm{BHL}}\approx 2.4\times 10^{-8}\,M_{\odot}$ yr-1. Inferring the comparable cloud capture accretion rate from Figure 13 yields a very similar rate. In general, we expect that tidal cloud capture may play a significant role in disc assembly, alongside BHL accretion. This may increase the rate of interstellar gas capture by a factor of order unity, and may be the subject of future numerical experiment.
4.1.3 Face-on accretion
A third mechanism for ISM accretion is face-on disc accretion, where an existing disc produces a large physical cross-section (Moeckel & Throop, 2009; Wijnen et al., 2017). In this case, the ram pressure between a pre-existing disc and the ISM mediates the accretion process. Face-on accretion may be significant when a star moves through a relatively high velocity medium. In this case, the disc shrinks as a result of deposition of low angular momentum material onto the disc surface (Wijnen et al., 2017). There is no steady state for this process unless material can be constantly replenished. This replenishment of outer disc material may be viscous, but then the quantity of material accreted is limited by the initial disc mass. Replenishment via BHL or cloud capture would keep the disc at a radius $R_{\mathrm{disc}}\sim R_{\mathrm{BHL}}$, and would therefore result in a similar accretion rate to BHL accretion (Moeckel & Throop, 2009). However, this may significantly enhance accretion rates when the BHL radius decreases; then $R_{\mathrm{disc}}>R_{\mathrm{BHL}}$. Face-on accretion would then temporarily dominate the overall rate, shrinking the disc to maintain a balance between shrinkage and replenishment. Overall, we may expect some moderate enhancement to the rate of capture and stellar accretion during short phases, although this would not be expected to greatly change the overall disc evolution.
4.2 Implications
@@
  • Disc sub-structures: A diverse range of substructures have been observed in protoplanetary discs (Andrews, 2020). We have demonstrated that the disc is undergoing constant perturbation by in-falling material. We would therefore expect some discs to exhibit structure due to this process (Kuffmeier et al., 2020). For example, by accreting gas from the ISM with a misaligned angular momentum vector we expect disc warping (Dullemond et al., 2022), which may also produce observable shadows (Marino et al., 2015; Krieger et al., 2024). Both external structures and disc warping are evident in some discs (Walsh et al., 2017), and it seems probable that they originate from late stage in-fall at least in some instances.
@@
  • Accretion outbursts and short accretion times: FU Orionis events are short accretion bursts that last $\sim 100$ years (Hartmann & Kenyon, 1996), named for the archetypal example (Herbig, 1966, 1977). In-fall events are among the earliest explanations for these events. There remain several other possible mechanisms such as stellar encounters (Bonnell & Bastien, 1992; Borchert et al., 2022) and thermal instability (Lin et al., 1985; Clarke & Syer, 1996; Lodato & Clarke, 2004), although extended nebulosity surrounding several stars undergoing outburst strongly supports the role of in-fall in at least some cases (Zurlo et al., 2024; Hales et al., 2024). Outbursting star-disc systems may be more broadly categorised as a subset of systems with short accretions time-scales (Manara et al., 2023; Almendros-Abad et al., 2024). Among such discs, a lifetime problem appears particularly pronounced for discs around some Herbig stars (Grant et al., 2023). The assumptions of our model mean that we cannot directly explain these short accretion time-scales; we impose this time-scale for our model. However, given a fixed time-scale $\tau_{\mathrm{acc}}$, discs with short $\tau_{\mathrm{acc}}$ may be episodically replenished in our model. Thus, to a degree the inferred accretion time-scale is decoupled from the expected frequency of observed discs, which no longer needs to be $\sim\tau_{\mathrm{acc}}/\tau_{\mathrm{disc}}$, for disc age $\tau_{\mathrm{disc}}$. Instead, the frequency of discs with a given $\tau_{\mathrm{acc}}$ would be dictated by the frequency of in-fall events that lead to rapid mass gain. Indeed, in our model a small number of discs with very short accretion time-scales ($\sim 0.01$ Myr) are retained for $\sim 3$ Myr (Figure 5). Whether the in-fall itself can also produce a short-lived dip in the accretion time-scale (i.e. an outburst) depends on the turbulence imparted onto the disc, and may depend on the rate or physics of interstellar gas accretion (as discussed above).
@@
  • Long-lived discs: Stochastic replenishment of discs represents a simple explanation for so-called ‘Peter Pan’ discs, which are (low mass) stars that are tens of Myr old but still exhibit a gas-rich accretion disc (Silverberg et al., 2020; Lee et al., 2020; Galli et al., 2021a). It is not clear if Peter Pan discs are exclusively around low mass stars, or whether finding discs around M dwarf stars is just down to sampling of the mass function. Speculatively, this phenomenon may be related to the fact that individual massive discs appear across a number of regions of different ages, even where the majority of discs have been depleted (Ansdell et al., 2020). Late stage in-fall seems a viable way to explain some of these long-lived discs. This further motivates deep observational programs to characterise the super-disc scale gas surrounding protoplanetary discs.
@@
  • Planet properties: Possibly the most exciting consequence of protoplanetary discs that are regulated by BHL accretion would be the potential consequences for planets. Multiple studies have shown that mature planet properties are dependent on stellar kinematics (Hamer & Schlaufman, 2019; Winter et al., 2020) Some of these differences could be related to covariance with stellar properties, such as age and metallicity (Miyazaki & Masuda, 2023). However, it is not clear that all correlations can be explained by stellar properties alone (Longmore et al., 2021; Winter & Alexander, 2021; Zink et al., 2023). If discs are constantly replenished by the ISM, stars that form at high velocity relative to the local rest frame should accrete less mass from the ISM after the dispersal of their formation cloud. This represents a direct theoretical link between planet formation and stellar kinematics.
5 Conclusions
In this work, we have presented a novel method for computing the evolution of rate at which protoplanetary discs capture material from their surroundings via Bondi-Hoyle-Lyttleton (BHL) accretion. To this end, we have adapted the excursion set formalism developed by Hopkins (2012) to track the local density and velocity evolution of stars throughout the lifetime of the protoplanetary disc. Our three most important conclusions from this experiment are:
  1. 1.
    Disc masses, radii and stellar accretion rates as a function of stellar mass and age can result directly from BHL accretion.
  2. 2.
  3. 3.
Finally, a key advantage of our method is that it can be easily adapted to semi-analytic models in order to include the role of BHL accretion on planet formation. This, as well as unbiased observational surveys to test the findings presented in this work, represent two urgent goals for future studies.
Acknowledgements
We thank Cathie Clarke, Alessandro Morbidelli and Sarah Jeffreson for useful discussion. AJW has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101104656. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (PROTOPLANETS, grant agreement No. 101002188).